# Consistency conditions for O-plane unsmearing from second-order perturbation theory

### v1

@@ 5 Conclusion
Confirming or disproving the consistency of DGKT-type scale-separated AdS compactifications with fully localized sources would be an important step in understanding questions of scale-separation in string theory. In this work, we have examined the perturbative O-plane unsmearing procedure for DGKT-type vacua at second order in the metric and dilaton perturbations. At this order the non-linearities of the supergravity equations of motion first manifest themselves. We investigated the possibility of additional non-linear constraints or inconsistencies, similar to linearization instabilities in general relativity, arising from integrating the second-order equations. We found that the singular near-source behavior of the first-order corrections leads to a consistent near-source behavior of the second-order corrections, provided one treats it as a generalized function, with an appropriate type of singular contribution on the source locus. This in turn allows for a consistent derivation of integral constraints for the first-order corrections. The system has enough degrees of freedom that these constraints do not lead to inconsistencies, but rather serve to fix the integration constants of the first-order corrections. We also found that O-plane intersections appear to have very benign behavior, at least at this order.

### v2

@@ 5 Conclusion
Confirming or disproving the consistency of DGKT-type scale-separated AdS compactifications with fully localized sources would be an important step in understanding questions of scale-separation in string theory. In this work, we have examined the perturbative O-plane unsmearing procedure for DGKT-type vacua at second order in the metric and dilaton perturbations. At this order the non-linearities of the supergravity equations of motion first manifest themselves. We investigated the possibility of additional non-linear constraints or inconsistencies, similar to linearization instabilities in general relativity, arising from integrating the second-order equations. We found that the singular near-source behavior of the first-order corrections leads to a consistent near-source behavior of the second-order corrections, provided one treats them as generalized functions, with an appropriate type of singular contribution on the source locus. This in turn allows for a consistent derivation of integral constraints for the first-order corrections. The system has enough degrees of freedom that these constraints do not lead to inconsistencies, but rather serve to fix the integration constants of the first-order corrections. We also found that O-plane intersections of the sort that can occur in DGKT-type constructions appear to have very benign behavior, at least at this order.